OpenStudy (anonymous):

how do i do this? Solve for x when 5(1-x)-2=3x-4????

OpenStudy (anonymous):

5-5x-2=3x-4 -5x+3=3x-4 -8x=-7 x=7/8

OpenStudy (anonymous):

x=7/8

OpenStudy (anonymous):

@ you got it .....................